Volume and cylindrical shell method Question 8 (10 points) Find the volume of the solid formed by rotating the region boundedQuestion 6 (10 points) Rotate the region bounded by y = 2x2 and y = x3 X about the X-axis
Answer #1

Here we will evaluate the volume of the given region using cylindrical shell method with proper limit.

8 Let S be the solid formed by rotating the region bounded by y= 0, y = In=0 and n=1 Itn² About the y-axis. We have y = ㅗ i+nL Here if n=0, then y = I 1 +0 and it n=1 then y = 14 thus (164)*= x2= y=14+ y = 1 and function 헌 Now the volume of the regio* [+(y)]?dy % -1) dy Sit [long -y] = 7 = * [ In 1-1 - kn + =ñ a In 2 - 2

Thus volume of the given region is

\pi \left( \ln 2- \frac{1}{2} \right).

LL 6 Ld s be the bounded by region $(^)= y = 2n2 and y= n = g(x), abood the x-axis Now, -y= 2n² = n3 = 2n² = n2 (0-2)=0 =n=02 NE [474 - n] dn Mso 2 大 45 곡] 1 4.25 2 7 1 大. ។ 2. 5 x 2.2 7 155 e V= 25 35 = 256 35

Thus the volume is

\pi \frac{256}{35}.